OpenStudy (aravindg):

How should I proceed to integrate this:

OpenStudy (shiraz14):

where's the question?

OpenStudy (fibonaccichick666):

He's getting there

OpenStudy (karategirl2002):

@hartnn @GabbyCakes @Mel98 @bringmethehorizon

OpenStudy (aravindg):

\[\large \int\limits \dfrac{\sec x \;dx}{\sqrt{\cos 2x}}\]

OpenStudy (karategirl2002):

@ElvenBlackout

OpenStudy (therealmeeeee):

@mathmale

OpenStudy (aravindg):

Why is this becoming a tagging party? :O

OpenStudy (fibonaccichick666):

I'd start by re-writing sec in sin and cos

OpenStudy (aravindg):

Let's just wait for someone to come up with a solution ;)

OpenStudy (anonymous):

Start by calling your alien friends, as they are a more developed race. especially more developed than an 8th grader like me. (AKA: I can't help you. I have no stinkin idea.)

OpenStudy (fibonaccichick666):

then use a double angle identity

OpenStudy (aravindg):

@FibonacciChick666 But which form should I express cos 2x?

OpenStudy (aravindg):

@BlazeBeat Calling all autobots :)

OpenStudy (fibonaccichick666):

the sin^2 one

OpenStudy (shiraz14):

I can't see the mathematical form of your integral - could you please use the equation function?

OpenStudy (anonymous):

@AravindG Autobots- Roll out!

OpenStudy (aravindg):

1-2sin^2x

OpenStudy (aravindg):

k lemme try it on paper.

OpenStudy (fibonaccichick666):

yea, then use u sub

OpenStudy (shiraz14):

OK ...

OpenStudy (aravindg):

@shiraz14 I used the equation editor above.

OpenStudy (shiraz14):

That's strange - all I see is code ...

OpenStudy (fibonaccichick666):

wait, that won't work... I think you'll run into a wall with the derivative

OpenStudy (anonymous):

ari- just wondering- did you watch any of transformers prime series?

OpenStudy (aravindg):

\[\dfrac{dx}{\cos x(\sqrt{1-2\sin^2 x})}\]

OpenStudy (aravindg):

Yeah thats what I landed on when I tried your method @FibonacciChick666

OpenStudy (aravindg):

@BlazeBeat All of them :P

OpenStudy (fibonaccichick666):

hmm have you had complex variables?

OpenStudy (karategirl2002):

@moonlight123

OpenStudy (shiraz14):

Ok, now I see the integral - its' a fault with my browser ... :p

OpenStudy (aravindg):

I don't think it's too complex for that. Then answer looks simple. I am missing out the method of attack. Wolf has the answer https://www.wolframalpha.com/input/?i=int+%28sec+x+dx%29%2F%28sqrt%7Bcos+2x%7D%29

OpenStudy (anonymous):

@AravindG I hated how it ended. optimus had to die D: but bumblebee got to kill megatron, died right after, but got to talk again after!!!

OpenStudy (fibonaccichick666):

ooh wait! just let u= cos2x

OpenStudy (fibonaccichick666):

then you can cancel stuff

OpenStudy (aravindg):

@BlazeBeat Let's discuss it in pm later. For now back to question :)

OpenStudy (anonymous):

@AravindG aye aye cap'n

OpenStudy (aravindg):

u=cos 2x That leaves me with du=-2sin 2xdx

OpenStudy (fibonaccichick666):

yes, so now, you need to get the same theta

OpenStudy (shiraz14):

Have you tried using u = cosx?

OpenStudy (shiraz14):

Then convert cos2x into 2 cos^2(x) - 1 = 2u^2-1

OpenStudy (aravindg):

But then how will you express dx in terms of u?

OpenStudy (fibonaccichick666):

du=-sinx dx then they cancel

OpenStudy (fibonaccichick666):

use sinx/sinx, that would probably work

OpenStudy (shiraz14):

sinx = square root [1-cos^2 (x)]

OpenStudy (aravindg):

hmm.. Sorry but I am not able to follow. Could u show the steps you are trying to tell me?

OpenStudy (fibonaccichick666):

yea, I think this is what he's getting at let u=cosx du=-sinx dx

OpenStudy (aravindg):

But there is no sin x on numerator?

OpenStudy (shiraz14):

\[dx = - \frac{ 1 }{ \sqrt{1-u^{2}} } du\]

OpenStudy (fibonaccichick666):

\[\int\frac{1}{sinx(2cos^2 x-1)^{.5}} dx\]

OpenStudy (fibonaccichick666):

brb class

OpenStudy (aravindg):

okay let me try that @shiraz14

OpenStudy (anonymous):

0_0 im in geometry and i just read this whole thing.....

OpenStudy (aravindg):

lol ^

OpenStudy (anonymous):

and i just kept getting more and more confused

OpenStudy (aravindg):

@shiraz14 By your method I reached a more complex looking integral.

OpenStudy (aravindg):

@hartnn , @AccessDenied any inputs?

OpenStudy (accessdenied):

I was going to try a Weierstrass sub just for funsies, just woke up so i'll see what happens there. lol

OpenStudy (anonymous):

@AravindG do u really need help or do mods at 99 just like posting really really difficult questions

OpenStudy (aravindg):

I can wait lol.

OpenStudy (aravindg):

@..... Not really. Mods are just normal people like with their own life and things to study. This qn came up when I was doing a practice session.

OpenStudy (aravindg):

I am sure the question isn' t as difficult as it seems to be. We are missing an important starting step. @ganeshie8 ?

OpenStudy (anonymous):

isnt as difficult as it seems to be..... it has more letters and symbols than numbers

OpenStudy (anonymous):

:'( what class is this

OpenStudy (shiraz14):

The integral may look complex at first, but it simplifies nicely into the following: \[-\int\limits_{?}^{?}[u ^{2}(2u ^{2}-1)(1-u ^{2})]^{-1/2} du\]

OpenStudy (anonymous):

0_0^

OpenStudy (aravindg):

Yeah I got something like that What do u do with it? So many terms in the denominator?

OpenStudy (aravindg):

@..... LOL

OpenStudy (anonymous):

its funny cause shiraz said it all looks like code to me.....same here shiraz

OpenStudy (shiraz14):

LOL ... that was my browser because it wasn't working ... :)

OpenStudy (accessdenied):

In short that didn't lead anywhere nice. [[ plain text work http://www.texpaste.com/n/1x7wft9n ]] will keep looking here.

OpenStudy (anonymous):

tanx/2 would help.

OpenStudy (aravindg):

@Abhishek619 Could you elaborate?

OpenStudy (anonymous):

plz do

OpenStudy (anonymous):